Tim and Judy mix two kinds of feed for pedigreed dogs. They wish to make 27 pounds of feed worth $0.26 per pound by mixing one kind worth $0.17 per pound with another worth $0.36 per pound. How many pounds of the cheaper kind should they use in the mix?
Answer by jorel1380(3719)

27x.26=7.02
m+n=27
.17m+.36n=7.02
17m+36n=702
17m+17n=459
19n=243
n=12.79
m=14.21
You need 14.21 lbs. of the cheaper feed for your mix..